# The Primordial Black Hole Abundance: The Broader, the Better

### v1

@@ Abstract
We show that the abundance of primordial black holes, if formed through the collapse of large fluctuations generated during inflation and unless the power spectrum of the curvature perturbation is very peaked, is always dominated by the broadest profile of the compaction function, even though statistically it is not the most frequent. The corresponding threshold is therefore 2/5. This result exacerbates the tension when combining the primordial black hole abundance with the signal seen by pulsar timing arrays and originated from gravitational waves induced by the same large primordial perturbations.
1 Introduction and Conclusions
Primordial Black Holes (PBHs) have emerged as one of the most interesting topics in cosmology in the last years (see Ref. [1] for a recent review). PBHs could explain both some of the signals from binary black hole mergers observed in gravitational wave detectors [2] and be an important component of the dark matter in the Universe.
One of the crucial parameter in PBHs physics is the relative abundance of PBHs with respect to the dark matter component. This quantity is not easy to calculate in the scenario in which PBHs are formed by the collapse of large fluctuations generated during inflation upon horizon re-entry. Indeed, the formation probability is very sensitive to tiny changes in the various ingredients, such as the critical threshold of collapse, the non-Gaussian nature of the fluctuations, the choice of the window function to define smoothed observables (see again Ref. [1] for a nice discussion on such issues), the nonlinear corrections entering in the calculation of the PBHs abundance from the nonlinear radiation transfer function and the determination of the true physical horizon crossing [3] and the appearance of an infinite tower of local, non-local and higher-derivative operators upon dealing with the nonlinear overdensity [4].
@@ 2 The compaction function
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t)e^{2\zeta(\bf x)}{\rm d}{\bf x}^{2},$ (2.1)
where $a(t)$ is the scale factor in terms of cosmic time. Cosmological perturbations may gravitationally collapse to form a PBH depending on the amplitude measured at the peak of the compaction function, defined as the mass excess compared to the background value within a given radius (see for instance Ref. [5])
$C({\bf x})=2\frac{M({\bf x},t)-M_{\rm b}({\bf x},t)}{R({\bf x},t)},$ (2.2)
@@
$R({\bf x},t)=a(t)re^{\zeta(\bf x)}$ (2.3)
with spherical coordinate radius $r$, centered around position ${\bf x}$ and evaluated at time $t$. The compaction directly measures the overabundance of mass in a region and is therefore better suited than the curvature perturbation for determining when an overdensity collapses into a PBH. Furthermore, the compaction has the advantage to be time-independent on superhorizon scales. It can be written in terms of the density contrast as
$C({\bf x})=\frac{2\rho_{\rm b}}{R({\bf x},t)}\int{\rm d}^{3}{\bf x}\,\delta({\bf x},t),$ (2.4)
@@
$\delta({\bf x},t)=-\frac{4}{9}\frac{1}{a^{2}H^{2}}e^{-2\zeta(\bf x)}\left(\nabla^{2}\zeta({\bf x})+\frac{1}{2}\left({\mathbf{\nabla}}\zeta({\bf x})\right)^{2}\right).$ (2.5)
@@
$C_{\zeta}(r)$ $=$ $-\frac{4}{3}r\zeta^{\prime}(r).$ (2.6)
@@
$C^{\prime}(r_{m})=C^{\prime}_{\zeta}(r_{m})\left[1-\frac{3}{4}C_{\zeta}(r_{m})\right]=0,$ (2.7)
@@
$C^{\prime\prime}(r_{m})=C^{\prime\prime}_{\zeta}(r_{m})\left[1-\frac{3}{4}C_{\zeta}(r_{m})\right],$ (2.8)
@@
$C^{\prime}(r_{m})=C^{\prime}_{\zeta}(r_{m})=0\,\,\,\,{\rm or}\,\,\,\,\zeta^{\prime}(r_{m})+r_{m}\zeta^{\prime\prime}(r_{m})=0.$ (2.9)
Consider now a family of compaction functions which have in common the same value of $r_{m}$, but a different curvature at the maximum parametrized by [6]
$q=-\frac{1}{4}\frac{r_{m}^{2}C^{\prime\prime}(r_{m})}{C(r_{m})},$ (2.10)
Numerically, it has been noticed that the critical threshold depends on the curvature at the peak of the compaction function [6, 7, 8]
$C_{c}(q)=\frac{4}{15}e^{-1/q}\frac{q^{1-5/2q}}{\Gamma(5/2q)-\Gamma(5/2q,1/q)},$ (2.11)
such that $C_{c}(q\ll 1)\simeq 2/5$ and $C_{c}(q\gg 1)\simeq 2/3$. We also notice that
$q=-\frac{1}{4}\frac{r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m})\left[1-\frac{3}{4}C_{\zeta}(r_{m})\right]}{C_{\zeta}(r_{m})\left[1-\frac{3}{8}C_{\zeta}(r_{m})\right]}\simeq-\frac{1}{4}\frac{r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m})}{C_{\zeta}(r_{m})}\left[1-\frac{3}{8}C_{\zeta}(r_{m})\right]\equiv q_{\zeta}\left[1-\frac{3}{8}C_{\zeta}(r_{m})\right].$ (2.12)
2.1 The average profile
@@
$\langle C_{\zeta}(r)\rangle_{\zeta(0)}=-\frac{4}{3}r\langle\zeta^{\prime}(r)\rangle_{\zeta(0)}=-\frac{4}{3}r\langle\zeta(r)\rangle_{\zeta(0)}^{\prime}=-\frac{4}{3}r\frac{\xi^{\prime}(r)}{\xi(0)}\zeta_{0},$ (2.13)
@@
$\xi(r)=\int\frac{{\rm d}k}{k}{\cal P}_{\zeta}(k)\frac{\sin kr}{kr}$ (2.14)
@@
$\xi^{\prime}(r_{m})+r_{m}\xi^{\prime\prime}(r_{m})=0.$ (2.15)
@@
$q=-\frac{1}{4}\frac{r_{m}^{2}\langle C^{\prime\prime}_{\zeta}(r_{m})\rangle_{\zeta(0)}}{\langle C_{\zeta}(r_{m})\rangle_{\zeta(0)}}\left[1-\frac{3}{8}\langle C_{\zeta}(r_{m})\rangle_{\zeta(0)}\right].$ (2.16)
@@
3 The relevance of Broadness: the Gaussian case
@@
$\sigma^{2}_{0}=\langle C_{\zeta}^{2}(r_{m})\rangle,\,\,\,\sigma^{2}_{1}=-\frac{1}{4}r_{m}^{2}\langle C_{\zeta}^{\prime\prime}(r_{m})C_{\zeta}(r_{m})\rangle,\,\,\,{\rm and}\,\,\,\sigma^{2}_{2}=\frac{1}{16}r_{m}^{4}\langle{C_{\zeta}^{\prime\prime}(r_{m})}^{2}\rangle.$ (3.1)
@@
$C_{\zeta}({\bf k},r)=\frac{4}{9}k^{2}r^{2}W(kr)\zeta({\bf k}),\,\,\,\,W(x)=3\frac{\sin x-x\cos x}{x^{3}},$ (3.2)
@@
$P\left[C(r_{m}),C^{\prime\prime}(r_{m})\right]{\rm d}C(r_{m}){\rm d}C^{\prime\prime}(r_{m})$ $=$ ${\cal P}\left[C_{\zeta}(r_{m}),C_{\zeta}^{\prime\prime}(r_{m})\right]{\rm d}C_{\zeta}(r_{m}){\rm d}C_{\zeta}^{\prime\prime}(r_{m})$ (3.3)
@@
$\Sigma$ $=$ $\left(\begin{array}[]{cc}\sigma^{2}_{2}&\sigma^{2}_{1}\\ \sigma^{2}_{1}&\sigma^{2}_{0}\end{array}\right).$ (3.6)
@@
$\gamma=\frac{\sigma_{1}^{2}}{\sigma_{2}\sigma_{0}},$ (3.7)
@@
3.1 The average of the curvature
The average curvature of the linear compaction function $C_{\zeta}$ can be computed by using the conditional probability to have a peak at $r_{m}$ [^2] In fact we use threshold statistics rather than peak statistics to elaborate our point. However, regions well above the corresponding square root of the variance are very likely local maxima [9].
$\langle q_{\zeta}\rangle=\int_{0}^{\infty}{\rm d}q_{\zeta}\,q_{\zeta}\,P[q|C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})],$ (3.8)
@@ 3.1 The average of the curvature
$P[q|C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]=\frac{\widetilde{{\cal P}}[q_{\zeta},C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]}{\widetilde{{\cal P}}[C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]}.$ (3.9)
@@
$P[q_{\zeta}|C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]$ $\simeq$ $\frac{\left(1-\gamma^{2}\right)^{1/2}\sigma_{2}C_{\zeta,c}(q_{\zeta})}{\sqrt{2\pi}\sigma_{0}\left[(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})^{2}+(1-\gamma^{2})\sigma_{2}^{2}/\sigma_{0}^{2}\right]^{1/2}}\cdot$ (3.10)
@@
$\langle q_{\zeta}\rangle\simeq\gamma\frac{\sigma_{2}}{\sigma_{0}}.$ (3.11)
3.2 The PBHs abundance
@@
$\beta=\int_{C_{c}(q)}^{\infty}{\rm d}C(r_{m})\int_{-\infty}^{0}{\rm d}C^{\prime\prime}(r_{m})P\left[C(r_{m}),C^{\prime\prime}(r_{m})\right]=\int_{0}^{\infty}{\rm d}q_{\zeta}\int_{C_{\zeta,c}(q_{\zeta})}^{\infty}{\rm d}C_{\zeta}(r_{m})\widetilde{{\cal P}}\left[C_{\zeta}(r_{m}),q_{\zeta}\right].$ (3.12)
@@
${\cal P}\left[-\frac{1}{4}r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m}),C_{\zeta}(r_{m})\right]$ $=$ $\frac{1}{2\pi}\frac{1}{\sigma_{2}\sigma_{0}\sqrt{1-\gamma^{2}}}{\rm exp}\left[-\frac{r_{m}^{4}{C_{\zeta}^{\prime\prime}(r_{m})}^{2}}{16\cdot 2\sigma_{2}^{2}}\right]\cdot$ (3.13)
@@
$\lim_{\gamma\rightarrow 1}{\cal P}\left[-\frac{1}{4}r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m}),C_{\zeta}(r_{m})\right]=\frac{1}{\sqrt{2\pi}}\frac{1}{\sigma_{2}\sigma_{0}}{\rm exp}\left(-\frac{r_{m}^{4}{C_{\zeta}^{\prime\prime}(r_{m})}^{2}}{16\cdot 2\sigma_{2}^{2}}\right)\delta_{D}\left(\frac{C_{\zeta}(r_{m})}{\sigma_{0}}+\frac{r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m})}{4\sigma_{2}}\right),$ (3.14)
@@
$q_{\zeta}=\frac{\sigma_{2}}{\sigma_{0}}=\frac{\sigma_{2}\sigma_{0}}{\sigma_{0}^{2}}=\frac{\sigma^{2}_{1}}{\sigma^{2}_{0}}=\langle q_{\zeta}\rangle,$ (3.15)
@@
$\beta$ $=$ $\int_{0}^{\infty}{\rm d}q_{\zeta}\int_{C_{\zeta,c}(q_{\zeta})}^{4/3}{\rm d}C_{\zeta}(r_{m}){\cal P}\left[-\frac{1}{4}r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m}),C_{\zeta}(r_{m})\right]$ (3.16)
@@
$C_{\zeta,c}(q_{\zeta})\simeq\frac{4}{3}\left(1-\sqrt{\frac{2-3C_{c}(q_{\zeta})}{2}}\right).$ (3.17)
@@
$C^{\rm peaked}_{\zeta,c}=C_{\zeta,c}(\langle q_{\zeta}\rangle).$ (3.18)
@@
$C^{2}_{\zeta,c}(q_{\zeta})\Big{|}_{\rm eff}=\left[(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})^{2}+(1-\gamma^{2})\sigma_{2}^{2}/\sigma_{0}^{2}\right]C^{2}_{\zeta,c}(q_{\zeta}).$ (3.20)
@@
$(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})C_{\zeta,c}(q_{\zeta})+\left[(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})^{2}+(1-\gamma^{2})\sigma_{2}^{2}/\sigma_{0}^{2}\right]\frac{{\rm d}C_{\zeta,c}(q_{\zeta})}{{\rm d}q_{\zeta}}=0.$ (3.21)
@@
$(q_{\zeta}-\sigma_{2}/\sigma_{0})\left[C_{\zeta,c}(q_{\zeta})+(q_{\zeta}-\sigma_{2}/\sigma_{0})\frac{{\rm d}C_{\zeta,c}(q_{\zeta})}{{\rm d}q_{\zeta}}\right]=0$ (3.22)
@@
$C^{2}_{\zeta,c}(q_{\zeta})\Big{|}_{\rm eff}=\left[q_{\zeta}^{2}+\sigma_{2}^{2}/\sigma_{0}^{2}\right]C^{2}_{\zeta,c}(q_{\zeta}).$ (3.23)
and the threshold $C_{\zeta,c}(q_{\zeta})$ is also minimized for small $q_{\zeta}$. There is in general a critical value of $q_{\zeta}$ for which the abundance is always dominated by the broad spectra. We can see this behaviour by plotting the curve $(q_{\zeta\;\rm{min}},\gamma)$ obtain from the Eq. (3.21), as shown in Fig. 2. As we start decreasing from $\gamma=1$ where the minimum is in $\sigma_{2}/\sigma_{0}$, also the value of $q_{\zeta,\;\rm{min}}$ decreases, up until a critical value $\gamma_{\rm crit}$. For values of $\gamma$ below this point the function $C_{\zeta,c}(q_{\zeta})|_{\rm{eff}}$ does not have a minimum, but is monotonically increasing with $q_{\zeta}$, hence the minimum lies at the boundary of the interval, i.e. $q_{\zeta}=0$. The transition is therefore very sharp after the critical value.
It is also possible to evaluate the position of this minimum for different values of the parameter $\sigma_{2}/\sigma_{0}$, as shown in Fig. 2. We can understand the behaviour because having larger values of this parameter the transition happens for larger values of $\gamma$, being easier to enter in the regime of Eq. (3.23), where $\sigma_{2}/\sigma_{0}$ dominates.
@@
$C_{\zeta,c}(q_{\zeta}\simeq 0)\simeq\frac{4}{3}\left(1-\sqrt{\frac{2-3\cdot 2/5}{2}}\right)\simeq 0.49.$ (3.24)
@@
4 The relevance of Broadness: the non-Gaussian case
As a matter of fact, the curvature perturbation generated in models producing large overdensities is typically non-Gaussian. Non-Gaussianity among the modes interested in the growth of the curvature perturbation is generated either by their self-interaction during the ultra slow-roll phase [10] or after Hubble radius exit when the curvature perturbation is sourced by a curvaton-like field [11, 12, 13].

### v2

@@ Abstract
We show that the abundance of primordial black holes, if formed through the collapse of large fluctuations generated during inflation and unless the power spectrum of the curvature perturbation is very peaked, is always dominated by the broadest profile of the compaction function, where the corresponding threshold is 2/5, even though statistically it is not the most frequent. This result exacerbates the tension when combining the primordial black hole abundance with the signal seen by pulsar timing arrays and originated from gravitational waves induced by the same large primordial perturbations.
I Introduction and Conclusions
Primordial Black Holes (PBHs) have emerged as one of the most interesting topics in cosmology in the last years (see Ref. LISACosmologyWorkingGroup:2023njw for a recent review). PBHs could explain both some of the signals from binary black hole mergers observed in gravitational wave detectors Franciolini:2021tla and be an important component of the dark matter in the Universe.
One of the crucial parameter in PBHs physics is the relative abundance of PBHs with respect to the dark matter component. This quantity is not easy to calculate in the scenario in which PBHs are formed by the collapse of large fluctuations generated during inflation upon horizon re-entry. Indeed, the formation probability is very sensitive to tiny changes in the various ingredients, such as the critical threshold of collapse, the non-Gaussian nature of the fluctuations, the choice of the window function to define smoothed observables (see again Ref. LISACosmologyWorkingGroup:2023njw for a nice discussion on such issues), the nonlinear corrections entering in the calculation of the PBHs abundance from the nonlinear radiation transfer function and the determination of the true physical horizon crossing DeLuca:2023tun and the appearance of an infinite tower of local, non-local and higher-derivative operators upon dealing with the nonlinear overdensity Franciolini:2023wun .
@@ II The compaction function
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t)e^{2\zeta(\bf x)}{\rm d}{\bf x}^{2},$ (1)
where $a(t)$ is the scale factor in terms of cosmic time. Cosmological perturbations may gravitationally collapse to form a PBH depending on the amplitude measured at the peak of the compaction function, defined as the mass excess compared to the background value within a given radius (see for instance Ref. Harada:2023ffo )
$C({\bf x})=2\frac{M({\bf x},t)-M_{\rm b}({\bf x},t)}{R({\bf x},t)},$ (2)
@@
$R({\bf x},t)=a(t)\tilde{r}e^{\zeta(\bf x)}$ (3)
with spherical coordinate radius $\tilde{r}$, centered around position ${\bf x}$ and evaluated at time $t$. The compaction directly measures the overabundance of mass in a region and is therefore better suited than the curvature perturbation for determining when an overdensity collapses into a PBH. Furthermore, the compaction has the advantage to be time-independent on superhorizon scales. It can be written in terms of the density contrast as
$C({\bf x})=\frac{2\rho_{\rm b}}{R({\bf x},t)}\int{\rm d}^{3}{\bf x}\,\delta({\bf x},t),$ (4)
@@
$\delta({\bf x},t)=-\frac{4}{9}\frac{1}{a^{2}H^{2}}e^{-2\zeta(\bf x)}\left(\nabla^{2}\zeta({\bf x})+\frac{1}{2}\left({\mathbf{\nabla}}\zeta({\bf x})\right)^{2}\right).$ (5)
@@
$C_{\zeta}(r)$ $=$ $-\frac{4}{3}r\zeta^{\prime}(r).$ (6)
@@
$C^{\prime}(r_{m})=C^{\prime}_{\zeta}(r_{m})\left[1-\frac{3}{4}C_{\zeta}(r_{m})\right]=0,$ (7)
@@
$C^{\prime\prime}(r_{m})=C^{\prime\prime}_{\zeta}(r_{m})\left[1-\frac{3}{4}C_{\zeta}(r_{m})\right],$ (8)
@@
$C^{\prime}(r_{m})=C^{\prime}_{\zeta}(r_{m})=0\,\,\,\,{\rm or}\,\,\,\,\zeta^{\prime}(r_{m})+r_{m}\zeta^{\prime\prime}(r_{m})=0.$ (9)
Consider now a family of compaction functions which have in common the same value of $r_{m}$, but a different curvature at the maximum parametrized by Escriva:2019phb
$q=-\frac{1}{4}\frac{r_{m}^{2}C^{\prime\prime}(r_{m})}{C(r_{m})},$ (10)
Numerically, it has been noticed that the critical threshold depends on the curvature at the peak of the compaction function Escriva:2019phb ; musco ; musconl
$C_{c}(q)=\frac{4}{15}e^{-1/q}\frac{q^{1-5/2q}}{\Gamma(5/2q)-\Gamma(5/2q,1/q)},$ (11)
such that $C_{c}(q\rightarrow 0)\simeq 2/5$ and $C_{c}(q\rightarrow\infty)\simeq 2/3$. We also notice that
$q=-\frac{1}{4}\frac{r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m})\left[1-\frac{3}{4}C_{\zeta}(r_{m})\right]}{C_{\zeta}(r_{m})\left[1-\frac{3}{8}C_{\zeta}(r_{m})\right]}\simeq-\frac{1}{4}\frac{r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m})}{C_{\zeta}(r_{m})}\left[1-\frac{3}{8}C_{\zeta}(r_{m})\right]\equiv q_{\zeta}\left[1-\frac{3}{8}C_{\zeta}(r_{m})\right].$ (12)
II.1 The average profile
@@
$\langle C_{\zeta}(r)\rangle_{\zeta(0)}=-\frac{4}{3}r\langle\zeta^{\prime}(r)\rangle_{\zeta(0)}=-\frac{4}{3}r\langle\zeta(r)\rangle_{\zeta(0)}^{\prime}=-\frac{4}{3}r\frac{\xi^{\prime}(r)}{\xi(0)}\zeta_{0},$ (13)
@@
$\xi(r)=\int\frac{{\rm d}k}{k}{\cal P}_{\zeta}(k)\frac{\sin kr}{kr}$ (14)
@@
$\xi^{\prime}(r_{m})+r_{m}\xi^{\prime\prime}(r_{m})=0.$ (15)
@@
$q=-\frac{1}{4}\frac{r_{m}^{2}\langle C^{\prime\prime}_{\zeta}(r_{m})\rangle_{\zeta(0)}}{\langle C_{\zeta}(r_{m})\rangle_{\zeta(0)}}\left[1-\frac{3}{8}\langle C_{\zeta}(r_{m})\rangle_{\zeta(0)}\right].$ (16)
@@
III The relevance of Broadness: the Gaussian case
@@
$\sigma^{2}_{0}=\langle C_{\zeta}^{2}(r_{m})\rangle,\,\,\,\sigma^{2}_{1}=-\frac{1}{4}r_{m}^{2}\langle C_{\zeta}^{\prime\prime}(r_{m})C_{\zeta}(r_{m})\rangle,\,\,\,{\rm and}\,\,\,\sigma^{2}_{2}=\frac{1}{16}r_{m}^{4}\langle{C_{\zeta}^{\prime\prime}(r_{m})}^{2}\rangle.$ (17)
@@
$C_{\zeta}({\bf k},r)=\frac{4}{9}k^{2}r^{2}W(kr)\zeta({\bf k}),\,\,\,\,W(x)=3\frac{\sin x-x\cos x}{x^{3}},$ (18)
@@
$P\left[C(r_{m}),C^{\prime\prime}(r_{m})\right]{\rm d}C(r_{m}){\rm d}C^{\prime\prime}(r_{m})$ $=$ ${\cal P}\left[C_{\zeta}(r_{m}),C_{\zeta}^{\prime\prime}(r_{m})\right]{\rm d}C_{\zeta}(r_{m}){\rm d}C_{\zeta}^{\prime\prime}(r_{m})$ (19)
@@
$\Sigma$ $=$ $\left(\begin{array}[]{cc}\sigma^{2}_{2}&\sigma^{2}_{1}\\ \sigma^{2}_{1}&\sigma^{2}_{0}\end{array}\right).$ (22)
@@
$\gamma=\frac{\sigma_{1}^{2}}{\sigma_{2}\sigma_{0}},$ (23)
@@
III.1 The average of the curvature
The average curvature of the linear compaction function $C_{\zeta}$ can be computed by using the conditional probability to have a peak at $r_{m}$ [^2] In fact we use threshold statistics rather than peak statistics to elaborate our point. However, regions well above the corresponding square root of the variance are very likely local maxima Hoffman:1985pu .
$\langle q_{\zeta}\rangle=\int_{0}^{\infty}{\rm d}q_{\zeta}\,q_{\zeta}\,P[q|C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})],$ (24)
@@ III.1 The average of the curvature
$P[q|C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]=\frac{\widetilde{{\cal P}}[q_{\zeta},C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]}{\widetilde{{\cal P}}[C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]}.$ (25)
@@
$P[q_{\zeta}|C_{\zeta}(r_{m})>C_{\zeta,c}(q_{\zeta})]$ $\simeq$ $\frac{\left(1-\gamma^{2}\right)^{1/2}\sigma_{2}C_{\zeta,c}(q_{\zeta})}{\sqrt{2\pi}\sigma_{0}\left[(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})^{2}+(1-\gamma^{2})\sigma_{2}^{2}/\sigma_{0}^{2}\right]^{1/2}}\cdot$ (26)
@@
$\langle q_{\zeta}\rangle\simeq\gamma\frac{\sigma_{2}}{\sigma_{0}}.$ (27)
Hence for very broad spectrum, $\gamma\rightarrow 0$, one has $\langle q_{\zeta}\rangle\rightarrow 0$.
III.2 The PBHs abundance
@@
$\beta=\int_{C_{c}(q)}^{\infty}{\rm d}C(r_{m})\int_{-\infty}^{0}{\rm d}C^{\prime\prime}(r_{m})P\left[C(r_{m}),C^{\prime\prime}(r_{m})\right]=\int_{0}^{\infty}{\rm d}q_{\zeta}\int_{C_{\zeta,c}(q_{\zeta})}^{\infty}{\rm d}C_{\zeta}(r_{m})\widetilde{{\cal P}}\left[C_{\zeta}(r_{m}),q_{\zeta}\right].$ (28)
@@
${\cal P}\left[-\frac{1}{4}r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m}),C_{\zeta}(r_{m})\right]$ $=$ $\frac{1}{2\pi}\frac{1}{\sigma_{2}\sigma_{0}\sqrt{1-\gamma^{2}}}{\rm exp}\left[-\frac{r_{m}^{4}{C_{\zeta}^{\prime\prime}(r_{m})}^{2}}{16\cdot 2\sigma_{2}^{2}}\right]\cdot$ (29)
@@
$\lim_{\gamma\rightarrow 1}{\cal P}\left[-\frac{1}{4}r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m}),C_{\zeta}(r_{m})\right]=\frac{1}{\sqrt{2\pi}}\frac{1}{\sigma_{2}\sigma_{0}}{\rm exp}\left(-\frac{r_{m}^{4}{C_{\zeta}^{\prime\prime}(r_{m})}^{2}}{16\cdot 2\sigma_{2}^{2}}\right)\delta_{D}\left(\frac{C_{\zeta}(r_{m})}{\sigma_{0}}+\frac{r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m})}{4\sigma_{2}}\right),$ (30)
@@
$q_{\zeta}=\frac{\sigma_{2}}{\sigma_{0}}=\frac{\sigma_{2}\sigma_{0}}{\sigma_{0}^{2}}=\frac{\sigma^{2}_{1}}{\sigma^{2}_{0}}=\langle q_{\zeta}\rangle,$ (31)
@@
$\beta$ $=$ $\int_{0}^{\infty}{\rm d}q_{\zeta}\int_{C_{\zeta,c}(q_{\zeta})}^{4/3}{\rm d}C_{\zeta}(r_{m}){\cal P}\left[-\frac{1}{4}r_{m}^{2}C_{\zeta}^{\prime\prime}(r_{m}),C_{\zeta}(r_{m})\right]$ (32)
@@
$C_{\zeta,c}(q_{\zeta})\simeq\frac{4}{3}\left(1-\sqrt{\frac{2-3C_{c}(q_{\zeta})}{2}}\right).$ (33)
@@
$C^{\rm peaked}_{\zeta,c}=C_{\zeta,c}(\langle q_{\zeta}\rangle).$ (34)
@@
{mdframed}
$C^{2}_{\zeta,c}(q_{\zeta})\Big{|}_{\rm eff}=\left[(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})^{2}+(1-\gamma^{2})\sigma_{2}^{2}/\sigma_{0}^{2}\right]C^{2}_{\zeta,c}(q_{\zeta}).$ (36)
@@
$(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})C_{\zeta,c}(q_{\zeta})+\left[(q_{\zeta}-\gamma\sigma_{2}/\sigma_{0})^{2}+(1-\gamma^{2})\sigma_{2}^{2}/\sigma_{0}^{2}\right]\frac{{\rm d}C_{\zeta,c}(q_{\zeta})}{{\rm d}q_{\zeta}}=0.$ (37)
@@
$(q_{\zeta}-\sigma_{2}/\sigma_{0})\left[C_{\zeta,c}(q_{\zeta})+(q_{\zeta}-\sigma_{2}/\sigma_{0})\frac{{\rm d}C_{\zeta,c}(q_{\zeta})}{{\rm d}q_{\zeta}}\right]=0$ (38)
@@
$C^{2}_{\zeta,c}(q_{\zeta})\Big{|}_{\rm eff}=\left[q_{\zeta}^{2}+\sigma_{2}^{2}/\sigma_{0}^{2}\right]C^{2}_{\zeta,c}(q_{\zeta}).$ (39)
and the threshold $C_{\zeta,c}(q_{\zeta})$ is also minimized for small $q_{\zeta}$. There is in general a critical value of $q_{\zeta}$ for which the abundance is always dominated by the broad spectra. We can see this behaviour by plotting the curve $(q_{\zeta\;\rm{min}},\gamma)$ obtain from the Eq. (37), as shown in Fig. 2. As we start decreasing from $\gamma=1$ where the minimum is in $\sigma_{2}/\sigma_{0}$, also the value of $q_{\zeta,\;\rm{min}}$ decreases, up until a critical value $\gamma_{\rm crit}$. For values of $\gamma$ below this point the function $C_{\zeta,c}(q_{\zeta})|_{\rm{eff}}$ does not have a minimum, but is monotonically increasing with $q_{\zeta}$, hence the minimum lies at the boundary of the interval, i.e. $q_{\zeta}=0$. The transition is therefore very sharp after the critical value.
It is also possible to evaluate the position of this minimum for different values of the parameter $\sigma_{2}/\sigma_{0}$, as shown in Fig. 2. We can understand the behaviour because having larger values of this parameter the transition happens for larger values of $\gamma$, being easier to enter in the regime of Eq. (39), where $\sigma_{2}/\sigma_{0}$ dominates.
@@
$C_{\zeta,c}(q_{\zeta}\simeq 0)\simeq\frac{4}{3}\left(1-\sqrt{\frac{2-3\cdot 2/5}{2}}\right)\simeq 0.49.$ (40)
@@
IV The relevance of Broadness: the non-Gaussian case
As a matter of fact, the curvature perturbation generated in models producing large overdensities is typically non-Gaussian. Non-Gaussianity among the modes interested in the growth of the curvature perturbation is generated either by their self-interaction during the ultra slow-roll phase cai or after Hubble radius exit when the curvature perturbation is sourced by a curvaton-like field curvaton ; curvaton_2 ; curvaton_3 .